OpenStudy (anonymous):

1 -x^2+2+3=0 what are the solutions? 2

OpenStudy (anonymous):

the easiest way to visualize this is multiply both sides by -1 to get the lead coefficient positive. then you will have \[x^2-2x-3\] then this is easily factored

OpenStudy (anonymous):

ax^2+bx+c=0 --> x1=(-b+(b^2-4ac)^1/2)/2a and x2=(-b-(b^2-4ac)^1/2)/2a

OpenStudy (gorv):

x^2-3x+x-3=0 x(x-3)+x-3=0 (x-3)(x+1)=0 x=-1 and x=3

OpenStudy (gorv):

@mandik1974

OpenStudy (anonymous):

;) thanks